# Edge Learning for 6G-enabled Internet of Things: A Comprehensive Survey of Vulnerabilities, Datasets, and Defenses

### v1

@@ VI-C Electricity theft detection
Electricity theft is a significant concern for utility systems as it causes high economic losses. Furthermore, the shift from controlled to smart devices may introduce highly sophisticated attacks that are harder to defend against. This requires AI-based defense mechanisms to mitigate the risks, which in turn requires Benin profile data to differentiate the pattern of attacks. Examples of such datasets include:
@@ VI-D Image classification
Image classification represents a powerful task for assessing the architectures and modern methodologies within the area of computer vision. Image classification is a core discipline concerned with understanding what an image looks like in its entirety, and the objective is to classify the given image with a specific category. Various datasets have been proposed to evaluate such security architectures and systems, including:
@@
  • Fashion-MNIST dataset [189]: Based on the broad success that the MNIST dataset has achieved, various works have attempted to take a similar approach. For example, Fashion-MNIST is designed to become a full-scale substitute for the original MNIST dataset when evaluating ML-based algorithms. The dataset shares the identical image size (70,000), the same data format (28x28 Grey-scale imagery), and the identical training (60,000 images) and testing (10,000 images) setup. However, instead of handwritten digits, Fashion-MNIST is made up of fashion items that fall under ten categories, each with 7,000 images per category. This dataset is also used for evaluating different edge-located security schemes such as in [164, 165, 166, 167].
  • FEMNIST dataset [190]: is a federated version of the MNIST dataset containing 341873 samples coming from more than 3300 writers as FL clients. With distinguishable writing styles, individual authors’ data is Non-Independent and Identically Distributed (Non-IID). Each client in the FEMNIST dataset has approximately 100 samples with 10 classes. This dataset has been used to evaluate edge-enabled FL-based security systems recently [171, 10].
  • CIFAR-10 and CIFAR-100 datasets [191]: represent labeled subsets from the 80 million tiny images dataset. The CIFAR-10 dataset is made up of 60,000 color images of 32x32 split into 10 classes, each with 6,000 images per class. It includes 50,000 training and 10,000 test images. The dataset is partitioned into 5 training batches and 1 test batch, each one containing 10000 images. The testing batch contains 1000 images randomly picked from every class. Each training batch contains the remaining images in a completely random order, although some of them may have more images within 1 class than any other. In addition, the training batches hold 5000 images from every class. The work by Xiang et al. [149] used this dataset to benchmark their proposed backdoor attacks defense mechanism.
  • ImageNet10 dataset [192]: is a hand-picked 10-class dataset derived from the ImageNet dataset, which provides high-resolution images divided into 1000 classes and formed only from the 320x320 core regions of the original images. The ImageNet10 classes were chosen to be optically distinct and to reflect natural objects. This dataset is used in different works for benchmarking along with other datasets such as MNIST and CIFAR. [178]
@@
  • Stanford Dogs dataset [193]: It was constructed from imagery and annotations from ImageNet for categorizing fine-grained images. The dataset features a collection of images representing 120 breeds of dogs worldwide. This dataset includes 20,580 images with 120 categories. This dataset can be used in conjunction with other datasets such as MNIST to benchmark edge learning systems [177].
@@ VI-D1 Facial recognition
Facial recognition refers to the identification or verification of a person’s identity based on their face. It catches, processes, and matches different features based on the details of the person’s face. Different datasets are used for such tasks, including:
  • AT&T dataset [195]: also known as the Database of Faces. The dataset consists of 400 grayscale (92×112) images of human faces taken from 40 different individuals, with a total of 10 images for each person. For certain individuals, the images were shot at varying times, differing in lighting and facial expressions. The work of Zhang et al. [175] evaluated a GAN-based poisoning attack on this dataset.
  • LFW dataset [196]: The Labeled Faces in the Wild is a collection of face photographs conceived to investigate the unconstrained face recognition task. It has over 13000 face images gathered from the web. Individual faces are labeled with their respective names. A total of 1680 of the pictured persons possess two or more separate photos. This dataset was also used for generating poisoned training samples in [161].
  • CASIA dataset [197]: collected in the period 2007-2010 by the Institute of Automation, Chinese Academy of Sciences (CASIA), including 725 face images. The dataset was designed for Near-infrared vs. Visible light (NIR-VIS) face recognition, with 640×480 resolutions, and with 1 to 22 VIS and 5 to 50 NIR face images per person. This dataset was also used in [161].
  • VGGFace2 dataset [198]: is composed of more than 3M images featuring 9131 famous people across a wide range of ethnic groups. All images are downloaded from Google and exhibit considerable variation regarding background, age, and lighting. Overall, the entire dataset is relatively balanced (40.7% for females and 59.3% for males). It includes 80-843 images for each individual. Fung et al. [147] used the VGGFace2 dataset for evaluating a poisoning Sybil’s defense mechanism.
  • CelebA dataset [199]: constructed by labeling selected images from the CelebFaces dataset, including 10000 celebrities, with 20 images for each one. And all photos are annotated with 40 face attributes and 5 key points. EdgeConnect is a two-stage adversarial scheme composed of an edge generator tracked by an image completion system, which can be used for scene generation. The scheme proposed in [200] is evaluated using the CelebA dataset.
@@ VI-E Time series classification
Time series classification involves ML techniques designed to analyze several classes of labeled time series data in order to forecast or categorize which class a new dataset falls into.
  • UCR time series classification archive [201]: was launched in 2002 and has since grown to become an extensive resource for the time series data mining field. While the original version of the archive contained 16 data sets, the archive has seen periodic updates over the years. A major expansion took place in 2015 when the archive expanded to 85 datasets. Various recent works on adversarial attacks and defenses against time series classification (TSC) problems have been proposed as discussed in [159].
@@ VI-F Human Activity Recognition
  • HARUS dataset [202]: gathered from a group of 30 volunteers with ages ranging between 19-48 years wearing a smartphone. The data considered about their activities are split into 2 classes (dynamic and static) with each having 3 activities. It includes a total of 10299 samples with 561 attributes. The work in [160] evaluated their proposed bilevel optimization framework, which is intended to compute optimal poisoning attacks on FL.
  • Parkinson dataset [203]: created to identify people with Parkinson’s disease. It consists of 5,875 observations from a series of biomedical readings of 31 people’s voices, 23 of whom have Parkinson’s disease. This dataset is used for evaluating the framework in [160].
TABLE XII: Attacks on federated ML algorithms.
@@ VI-H Location awareness
  • Places database [207]: is designed to provide AI systems with high-level visual perception training, including background, recognition, and prediction tasks. It includes over 10M images covering over 400 specific scenes, with anywhere from 5000 to 30000 training images per class. In addition, this dataset is used for the evaluation of the EdgeConnect system discussed above [200].
@@ VI-I Text classification
Manual text classification involves analyzing the text material and making classifications accordingly, generally conducted by an annotator. Considering the fact that text is among the widest kinds of unstructured data, it is difficult and costly to parse, comprehend, arrange, and classify textual data. The purpose of automated text classification is to categorize documents into pre-specified classes, usually using Natural Language Processing (NLP) and ML algorithms. Diverse parties have provided datasets for facilitating such tasks, like:
  • DBPedia datasets [208]: was conceived as a way to mine organized content from Wikipedia’s entries. It consists of an archive of information that provides hierarchical and classification data for more than 4.23M instances covering 685 classes. Variants of this data collection are a widely used reference for evaluating NLP/text classification projects, giving a valuable baseline for multi-class/multi-label hierarchical text classification. For instance, the DBPedia-14 version which consists of 560000 instances from 14 ontological categories (including film and animal classes), was employed in [167] where the authors studied the impact of network-level adversaries on FL training and dropping attacks amplification.
@@
TABLE XIII: Machine learning vulnerabilities.
@@
Black-box attack: the configuration of the target ML models is unavailable to adversaries
Black-box adversarial attack
@@
White-box attacks require full and complete knowledge of the model being targeted, including its training method, architecture and parameter values
@@
- There are two distinct attack scenarios black-box and white-box attacks
@@
Falsifying machine learning training data to generate unwanted results in a decentralized mode. The adversary’s goal is to modify the parameters learned and manipulate the training data on their own device
@@
- Black-box attacks, the attacker is assumed to know the learning algorithm and feature set, and collect a substitute data set but has no knowledge of the training set as well as the trained parameters
@@
By monitoring more rounds of the FL system, the adversary can drop the network traffic from selected clients
@@
VI-J Highlights of Peer-to-Peer Features in Edge Learning across Various Dataset Types
@@
  • Multi-purpose security: IoT devices gather data on potential security threats, learning from it and sharing insights or models with other devices, while maintaining the privacy of raw data.
  • Attack classification: In network infrastructures, devices collect and process data on different types of network attacks, updating learning models based on individual experiences, and sharing these models with other devices.
@@
  • Human Activity Recognition (HAR): Wearable devices like smartwatches learn to recognize human activity from user data, and share their learning models, maintaining the privacy of the raw data.
  • Sentiment classification: Devices (e.g., smartphones) classify sentiment on local data (such as social media posts, and messages) and share the learned models while keeping the actual data private.
@@ VII Machine learning vulnerabilities
Machine learning systems are vulnerable to attack, which places limitations on the application of machine learning, especially in 6G–IoT Networks [215]. There are six attacks on federated ML algorithms, including, FL model extraction attack, FL evasion attack, FL inversion attack, FL inference attack, FL poisoning attack, and FL drooping attack, as presented in Table XII. For the purpose of visualizing some of these threats, Figure 7 illustrates various attacks against different ML paradigms including centralized, distributed, and decentralized learning during different model phases, including training, post-training, and inference phases. And while there are different strategies for targeting AI models, for instance in centralized learning by poisoning the data on which the model trains directly, while in decentralized learning with poisoned updates, the result remains the same: a compromised model. In this section, we provide a comprehensive classification of threats against ML, as presented in Figure 8. In addition, a thoughtful and comprehensive review of ML vulnerabilities based on the provided classification, and related to three main aspects, namely, the knowledge already acquired by the attacker, the type of attack employed, and the final objective, which is provided in Table XIII.
@@ VII-A Pre/Post-deployment attacks
Table XII lists the attacks that can be launched on federated machine learning (FL) algorithms along with their goals and descriptions. These attacks can be categorized based on the phase at which they occur - pre-deployment or post-deployment. Pre-deployment attacks target the aggregate model itself before it is deployed to honest clients. Post-deployment attacks occur after the aggregate model has been deployed to honest clients. The following six attacks compromise the privacy of the private data of honest clients [160, 216, 53, 217]:
@@ VII-B Backdoor attacks
Backdoor attacks involve the manipulation of training data or models to trigger specific behaviors in the model during the testing phase. Different types of backdoor attacks are presented, including RNN backdoor attacks, imperceptible backdoor pattern attacks, distributed backdoor attacks, pixel-space backdoor attacks, and clean-label backdoor attacks.
@@ VII-B1 RNN backdoor attacks
Backdoor attacks in deep neural networks are based on two assumptions, namely, the adversary takes control of the global model training operation and the second assumption is that the adversary only has control over particular training data. Chen et al. [146] proposed a mitigating system against backdoor attacks in the text classification based on Backdoor Keyword Identification. The proposed system can identify the poisoning patterns in the training dataset with no trust data and no knowledge of the backdoor release. The following text classification datasets are used in the performance evaluation with the LSTM models, Reuters-21578 dataset, 20 newsgroups, DBpedia ontology, and IMDB. The results show good results in terms of identification precision and recall of poisoning samples. To remove poisoning data, the authors proposed the following formula to filter the keywords in a dictionary and detect the suspicious keyword, which is most susceptible to be a backdoor keyword:
@@ VII-B2 Imperceptible backdoor pattern
In image classification, the attacker poisons the training set with a small ensemble of pictures sourced from a source class (or classes), incorporated with a backdoor feature, and then tagged to a target class. Xiang et al. [149] proposed a mitigation technique based on imperceptible backdoor patterns. The proposed technique can identify if the training set is poisoned and precisely recognizes the target class and training images in which the backdoor pattern is integrated. At the end of the process, the proposed technique performs reverse engineering to provide an estimate of this backdoor pattern used by the attacker. The CIFAR-10 dataset is used as a benchmark experiment, and the results show a reduced rate compared to state-of-the-art to no more than 4.9%.
The poisoned dataset used for training the victim classifier consists of a combination of the ${Dataset}_{backdoor}$ and the benign ${Dataset}_{train}$. The ${Dataset}_{backdoor}$ is a set of examples that have been intentionally modified with a backdoor pattern to trick the classifier into predicting a specific target class when presented with an image from a particular source class. The ${Dataset}_{backdoor}$ is defined as follows:
@@
$f\left(x;;{pattern}^{*}\right)$ is the embedding function that maps an image $x$ to a feature vector, and ${pattern}^{*}$ is the backdoor pattern that is added to the image. The resulting image $f\left(x;;{pattern}^{*}\right)$ is then labeled as ${Class}^{*}$, which is the target class that the backdoor is designed to trigger. The source class(es) $P_{{Source}^{*}}$ determines which images the backdoor is applied to. If $P_{{Source}^{*}}$ contains multiple classes, the backdoor can be triggered by images from any of those classes. The poisoned dataset is used to train the victim classifier, which learns to classify images from both the benign dataset and the poisoned dataset. However, since the poisoned dataset contains examples with backdoors, the victim classifier can be manipulated to produce incorrect predictions on images that contain the backdoor pattern.

### v2

@@ VI-C Electricity theft detection
Electricity theft is a significant concern for utility systems as it causes high economic losses. Furthermore, shifting from control to smart devices may introduce highly sophisticated attacks that are harder to defend against. This requires AI-based defense mechanisms to mitigate the risks, which in turn requires Benin profile data to differentiate the pattern of attacks. Examples of such datasets include:
@@ VI-D Image classification
Image classification represents a mighty task for assessing the architectures and modern methodologies within computer vision. Image classification is a core discipline concerned with understanding what an image looks like in its entirety, and the objective is to classify the given image with a specific category. Various datasets have been proposed to evaluate such security architectures and systems, including:
@@
  • Fashion-MNIST dataset [209]: Based on the broad success that the MNIST dataset has achieved, various works have attempted to take a similar approach. For example, Fashion-MNIST is designed to become a full-scale substitute for the original MNIST dataset when evaluating ML-based algorithms. The dataset shares the identical image size (70,000), the same data format (28x28 Grey-scale imagery), and the identical training (60,000 images) and testing (10,000 images) setup. However, instead of handwritten digits, Fashion-MNIST comprises fashion items that fall under ten categories, each with 7,000 images per category. This dataset is also used to evaluate edge-located security schemes such as in [185, 186, 187, 188].
  • FEMNIST dataset [210]: is a federated version of the MNIST dataset containing 341873 samples from over 3300 writers as FL clients. With distinguishable writing styles, individual authors’ data is Non-Independent and Identically Distributed (Non-IID). Each client in the FEMNIST dataset has approximately 100 samples with 10 classes. This dataset has been used to evaluate edge-enabled FL-based security systems recently [191, 13].
  • CIFAR-10 and CIFAR-100 datasets [211]: represent labeled subsets from the 80 million tiny images dataset. The CIFAR-10 dataset comprises 60,000 color images of 32x32 split into 10 classes, each with 6,000 images per class. It includes 50,000 training and 10,000 test images. The dataset is partitioned into 5 training batches and 1 test batch, each containing 10000 images. The testing batch contains 1000 images randomly picked from every class. Each training batch contains the remaining images randomly, although some may have more images within 1 class than others. In addition, the training batches hold 5000 images from every class. The work by Xiang et al. [170] used this dataset to benchmark their proposed backdoor attacks defense mechanism.
  • ImageNet10 dataset [212]: is a hand-picked 10-class dataset derived from the ImageNet dataset, which provides high-resolution images divided into 1000 classes and formed only from the 320x320 core regions of the original images. The ImageNet10 classes were chosen to be optically distinct and to reflect natural objects. This dataset is used in different works for benchmarking and other datasets such as MNIST and CIFAR. [198]
@@
  • Stanford Dogs dataset [213]: It was constructed from imagery and annotations from ImageNet for categorizing fine-grained images. The dataset features a collection of images representing 120 breeds of dogs worldwide. This dataset includes 20,580 images with 120 categories. This dataset can be used with other datasets such as MNIST to benchmark edge learning systems [197].
@@ VI-D1 Facial recognition
Facial recognition refers to identifying or verifying a person’s identity based on their face. It catches, processes, and matches different features based on the details of the person’s face. Different datasets are used for such tasks, including:
  • AT&T dataset [215]: also known as the Database of Faces. The dataset consists of 400 grayscale (92×112) images of human faces taken from 40 individuals, with 10 images for each person. For certain individuals, the images were shot at varying times, differing in lighting and facial expressions. The work of Zhang et al. [195] evaluated a GAN-based poisoning attack on this dataset.
  • LFW dataset [216]: The Labeled Faces in the Wild is a collection of face photographs conceived to investigate the unconstrained face recognition task. It has over 13000 face images gathered from the web. Individual faces are labeled with their respective names. 1680 of the pictured persons possess two or more separate photos. This dataset was also used for generating poisoned training samples in [182].
  • CASIA dataset [217]: collected in the period 2007-2010 by the Institute of Automation, Chinese Academy of Sciences (CASIA), including 725 face images. The dataset was designed for Near-infrared vs. Visible light (NIR-VIS) face recognition, with 640×480 resolutions and with 1 to 22 VIS and 5 to 50 NIR face images per person. This dataset was also used in [182].
  • VGGFace2 dataset [218]: comprises more than 3M images featuring 9131 famous people across a wide range of ethnic groups. All images are downloaded from Google and exhibit considerable background, age, and lighting variation. Overall, the entire dataset is relatively balanced (40.7% for females and 59.3% for males). It includes 80-843 images for each individual. Fung et al. [168] used the VGGFace2 dataset for evaluating a poisoning Sybil’s defense mechanism.
  • CelebA dataset [219]: constructed by labeling selected images from the CelebFaces dataset, including 10000 celebrities, with 20 images for each one. And all photos are annotated with 40 face attributes and five key points. EdgeConnect is a two-stage adversarial scheme composed of an edge generator tracked by an image completion system, which can be used for scene generation. The scheme proposed in [220] is evaluated using the CelebA dataset.
@@ VI-E Time series classification
Time series classification involves ML techniques designed to analyze several classes of labeled time series data to forecast or categorize which class a new dataset falls into.
  • UCR time series classification archive [221]: was launched in 2002 and has since become an extensive resource for the time series data mining field. While the original version of the archive contained 16 data sets, the archive has seen periodic updates over the years. A major expansion occurred in 2015 when the archive expanded to 85 datasets. Recent works on adversarial attacks and defenses against time series classification (TSC) problems have been proposed as discussed in [180].
@@ VI-F Human Activity Recognition
  • HARUS dataset [222]: gathered from a group of 30 volunteers with ages ranging between 19-48 years wearing a smartphone. The data about their activities are split into two classes (dynamic and static) with each having 3 activities. It includes a total of 10299 samples with 561 attributes. The work in [181] evaluated their proposed bilevel optimization framework, which is intended to compute optimal poisoning attacks on FL.
  • Parkinson dataset [223]: created to identify people with Parkinson’s disease. It consists of 5,875 observations from a series of biomedical readings of 31 people’s voices, 23 of whom have Parkinson’s disease. This dataset evaluates the framework in [181].
TABLE XIII: Attacks on federated ML algorithms.
@@ VI-H Location awareness
  • Places database [227]: is designed to provide AI systems with high-level visual perception training, including background, recognition, and prediction tasks. It includes over 10M images covering over 400 specific scenes, ranging from 5000 to 30000 training images per class. In addition, this dataset is used to evaluate the EdgeConnect system discussed above [220].
@@ VI-I Text classification
Manual text classification involves analyzing the text material and making classifications accordingly, generally conducted by an annotator. Because text is among the widest kinds of unstructured data, it is difficult and costly to parse, comprehend, arrange, and classify textual data. Automated text classification categorizes documents into pre-specified classes, usually using Natural Language Processing (NLP) and ML algorithms. Diverse parties have provided datasets for facilitating such tasks, like:
  • DBPedia datasets [228]: was conceived to mine organized content from Wikipedia’s entries. It consists of an archive of information that provides hierarchical and classification data for more than 4.23M instances covering 685 classes. Variants of this data collection are a widely used reference for evaluating NLP/text classification projects, giving a valuable baseline for multi-class/multi-label hierarchical text classification. For instance, the DBPedia-14 version which consists of 560000 instances from 14 ontological categories (including film and animal classes), was employed in [188] where the authors studied the impact of network-level adversaries on FL training and dropping attacks amplification.
@@
TABLE XIV: Machine learning vulnerabilities.
@@
blue-box attack: the configuration of the target ML models is unavailable to adversaries
blue-box adversarial attack
@@
White-box attacks require complete knowledge of the model being targeted, including its training method, architecture, and parameter values
@@
- There are two distinct attack scenarios blue-box and white-box attacks
@@
Falsifying machine learning training data to generate unwanted results in a decentralized mode. The adversary’s goal is to modify the parameters learned and manipulate the training data on their device
@@
- blue-box attacks, the attacker is assumed to know the learning algorithm and feature set, and collect a substitute data set but does not know the training set as well as the trained parameters
@@
By monitoring more rounds of the FL system; the adversary can drop the network traffic from selected clients
@@
VI-J Application of Dataset Types in 6G IoT Security
The categories of datasets, when employed in machine learning models, substantially bolster security in edge learning environments. IoT-based multi-purpose security datasets enable the detection of unusual cyber activities in real-time, safeguarding IoT networks [235]. Attack classification data can inform the creation of sophisticated intrusion detection systems to protect against known attack vectors [236]. Electricity theft detection through anomalous usage patterns ensures the integrity of the power grid, which indirectly benefits the security of all connected devices. Image classification allows for local object detection and facial recognition, which is crucial for many security applications in edge devices [237]. Time series classification can pinpoint sudden changes in network traffic indicative of a cyber attack [238]. Human Activity Recognition can identify abnormal or unauthorized activities in secure areas [239]. Sentiment classification aids in spotting harmful or threatening digital communications [240]. Location awareness is crucial in verifying user authenticity based on their usual geographic patterns, while text classification can help identify spam or phishing attempts [241].
VI-K Highlights of Peer-to-Peer Features in Edge Learning across Various Dataset Types
@@
  • Multi-purpose security: IoT devices gather data on potential security threats, learning from it and sharing insights or models with other devices while maintaining the privacy of raw data.
  • Attack classification: In network infrastructures, devices collect and process data on different network attacks, updating learning models based on individual experiences and sharing these models with other devices.
@@
  • Human Activity Recognition (HAR): Wearable devices like smartwatches learn to recognize human activity from user data and share their learning models, maintaining the privacy of the raw data.
  • Sentiment classification: Devices (e.g., smartphones) classify sentiment on local data (such as social media posts and messages) and share the learned models while keeping the actual data private.
@@ VII Machine learning vulnerabilities
Machine learning systems are vulnerable to attack, which limits the application of machine learning, especially in 6G–IoT Networks [242]. There are six attacks on federated ML algorithms, including FL model extraction attack, FL evasion attack, FL inversion attack, FL inference attack, FL poisoning attack, and FL drooping attack, as presented in Table XIII. To visualize some of these threats, Figure 9 illustrates various attacks against different ML paradigms including centralized, distributed, and decentralized learning during different model phases, including training, post-training, and inference phases. And while there are different strategies for targeting AI models, for instance in centralized learning by poisoning the data on which the model trains directly, in decentralized learning with poisoned updates, the result remains the same: a compromised model. This section provides a comprehensive classification of threats against ML, as presented in Figure 10. In addition, a thoughtful and comprehensive review of ML vulnerabilities based on the provided classification and related to three main aspects, namely, the knowledge already acquired by the attacker, the type of attack employed, and the final objective, which is provided in Table XIV.
@@ VII-A Pre/Post-deployment attacks
Table XIII lists the attacks that can be launched on federated machine learning (FL) algorithms along with their goals and descriptions. These attacks can be categorized based on the phase they occur - pre-deployment or post-deployment. Pre-deployment attacks target the aggregate model before being deployed to honest clients. Post-deployment attacks occur after the aggregate model has been deployed to honest clients. The following six attacks compromise the privacy of the private data of honest clients [181, 243, 120, 244]:
@@ VII-B Backdoor attacks
Backdoor attacks involve manipulating training data or models to trigger specific behaviors in the model during the testing phase. Different backdoor attacks are presented, including RNN backdoor attacks, imperceptible backdoor pattern attacks, distributed backdoor attacks, pixel-space backdoor attacks, and clean-label backdoor attacks.
@@ VII-B1 RNN backdoor attacks
Backdoor attacks in deep neural networks are based on two assumptions, namely, the adversary takes control of the global model training operation and the second assumption is that the adversary only has control over particular training data. RNN backdoor attacks on 6G could secretly embed malicious behaviors in network models, potentially compromising data integrity and user privacy. As 6G relies heavily on AI-driven solutions for efficient network management and data processing, such attacks can degrade network performance and security. Consequently, they pose risks to the expected ultra-reliable, low-latency communication features of 6G. Chen et al. [167] proposed a mitigating system against backdoor attacks in the text classification based on Backdoor Keyword Identification. The proposed system can identify the poisoning patterns in the training dataset with no trust data and no knowledge of the backdoor release. The following text classification datasets are used in the performance evaluation with the LSTM models, Reuters-21578 dataset, 20 newsgroups, DBpedia ontology, and IMDB. The results show good results regarding identification precision and recall of poisoning samples. To remove poisoning data, the authors proposed the following formula to filter the keywords in a dictionary and detect the suspicious keyword, which is most susceptible to be a backdoor keyword:
@@ VII-B2 Imperceptible backdoor pattern
In image classification, the attacker poisons the training set with a small ensemble of pictures sourced from a source class (or classes), incorporated with a backdoor feature, and then tagged to a target class. The imperceptible backdoor pattern in 6G can subtly introduce vulnerabilities, enabling unauthorized access or control. This compromises the integrity and security of ultra-reliable, high-speed 6G networks. As a result, it jeopardizes user privacy, data confidentiality, and the dependability of critical applications. Xiang et al. [170] proposed a mitigation technique based on imperceptible backdoor patterns. The proposed technique can identify if the training set is poisoned and precisely recognizes the target class and training images in which the backdoor pattern is integrated. At the end of the process, the proposed technique performs reverse engineering to provide an estimate of this backdoor pattern used by the attacker. The CIFAR-10 dataset is used as a benchmark experiment, and the results show a reduced rate compared to state-of-the-art to no more than 4.9%.
The poisoned dataset used for training the victim classifier consists of a combination of the ${Dataset}_{backdoor}$ and the benign ${Dataset}_{train}$. The ${Dataset}_{backdoor}$ is a set of examples intentionally modified with a backdoor pattern to trick the classifier into predicting a specific target class when presented with an image from a particular source class. The ${Dataset}_{backdoor}$ is defined as follows:
@@
$f\left(x;;{pattern}^{*}\right)$ is the embedding function that maps an image $x$ to a feature vector, and ${pattern}^{*}$ is the backdoor pattern that is added to the image. The resulting image $f\left(x;;{pattern}^{*}\right)$ is then labeled as ${Class}^{*}$, which is the target class that the backdoor is designed to trigger. The source class(es) $P_{{Source}^{*}}$ determines which images the backdoor is applied to. If $P_{{Source}^{*}}$ contains multiple classes, the backdoor can be triggered by images from any of those classes. The poisoned dataset is used to train the victim classifier, which learns to classify images from both benign and poisoned datasets. However, since the poisoned dataset contains examples with backdoors, the victim classifier can be manipulated to produce incorrect predictions on images that contain the backdoor pattern.